Find the solution of the logarithmic equation log(3x+2) = 3 in terms of logarithms, or correct to 4 decimal places. x = ?
Answer by nerdybill(7384)

log(3x+2) = 3
(3x+2) = 10^3
3x+2 = 1000
3x = 998
x = 332.6667